The optical products division of Panasonic is planning a $3.5 million building expansion for manufacturing its powerful Lumix DMC digital zoom camera. If the company uses an interest rate of 16% per year, compounded quarterly for all new investments, what is the uniform amount per quarter the company must make in order to recover its investment in 6 years?
Answer #1

Quarterly interest rate=16%/4=4%

Investment=$3.5 million

n=6*4=24

PW=A*(1-(1+r)^-n)/r

Or, 3500000=A*(1-(1+4%)^-24)/4%

Or, 3500000=A*(1-1.04^-24)/0.04

Or, 3500000=A*(1-0.3901)/0.04

Or, A=3500000*0.04/0.6099

Or, A=$ 229553.91

Hence uniform amount =$229553.91